\begin{thebibliography}{22}
\expandafter\ifx\csname natexlab\endcsname\relax\def\natexlab#1{#1}\fi
\expandafter\ifx\csname bibnamefont\endcsname\relax
  \def\bibnamefont#1{#1}\fi
\expandafter\ifx\csname bibfnamefont\endcsname\relax
  \def\bibfnamefont#1{#1}\fi
\expandafter\ifx\csname citenamefont\endcsname\relax
  \def\citenamefont#1{#1}\fi
\expandafter\ifx\csname url\endcsname\relax
  \def\url#1{\texttt{#1}}\fi
\expandafter\ifx\csname urlprefix\endcsname\relax\def\urlprefix{URL }\fi
\providecommand{\bibinfo}[2]{#2}
\providecommand{\eprint}[2][]{\url{#2}}

\bibitem[{\citenamefont{Gualtieri et~al.}(1982)\citenamefont{Gualtieri,
  Kincaid, and Morrison}}]{GKM82}
\bibinfo{author}{\bibfnamefont{J.~A.} \bibnamefont{Gualtieri}},
  \bibinfo{author}{\bibfnamefont{J.~M.} \bibnamefont{Kincaid}},
  \bibnamefont{and} \bibinfo{author}{\bibfnamefont{G.}~\bibnamefont{Morrison}},
  \bibinfo{journal}{J. Chem. Phys.} \textbf{\bibinfo{volume}{77}},
  \bibinfo{pages}{521} (\bibinfo{year}{1982}).

\bibitem[{\citenamefont{Sollich et~al.}(2001)\citenamefont{Sollich, Warren, and
  Cates}}]{SWC01}
\bibinfo{author}{\bibfnamefont{P.}~\bibnamefont{Sollich}},
  \bibinfo{author}{\bibfnamefont{P.~B.} \bibnamefont{Warren}},
  \bibnamefont{and} \bibinfo{author}{\bibfnamefont{M.~E.} \bibnamefont{Cates}},
  \bibinfo{journal}{Adv. Chem. Phys.} \textbf{\bibinfo{volume}{116}},
  \bibinfo{pages}{265} (\bibinfo{year}{2001}),
  \bibinfo{note}{cond-mat/0003084}.

\bibitem[{\citenamefont{Sollich}(2002)}]{S02}
\bibinfo{author}{\bibfnamefont{P.}~\bibnamefont{Sollich}}, \bibinfo{journal}{J.
  Phys.: Cond. Matt.} \textbf{\bibinfo{volume}{14}}, \bibinfo{pages}{R79}
  (\bibinfo{year}{2002}).

\bibitem[{\citenamefont{Blaak}(1998)}]{B98}
\bibinfo{author}{\bibfnamefont{R.}~\bibnamefont{Blaak}}, \bibinfo{journal}{Mol.
  Phys.} \textbf{\bibinfo{volume}{95}}, \bibinfo{pages}{695}
  (\bibinfo{year}{1998}).

\bibitem[{\citenamefont{Bartlett}(1999)}]{B99b}
\bibinfo{author}{\bibfnamefont{P.}~\bibnamefont{Bartlett}},
  \bibinfo{journal}{Mol. Phys.} \textbf{\bibinfo{volume}{97}},
  \bibinfo{pages}{685} (\bibinfo{year}{1999}).

\bibitem[{\citenamefont{Ogarko and Luding}(2012)}]{OL12}
\bibinfo{author}{\bibfnamefont{V.}~\bibnamefont{Ogarko}} \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{S.}~\bibnamefont{Luding}},
  \bibinfo{journal}{J. Chem. Phys.} \textbf{\bibinfo{volume}{136}},
  \bibinfo{pages}{124508} (\bibinfo{year}{2012}).

\bibitem[{\citenamefont{Santos et~al.}(1999)\citenamefont{Santos, Yuste, and
  {L\'opez de Haro}}}]{SYH99}
\bibinfo{author}{\bibfnamefont{A.}~\bibnamefont{Santos}},
  \bibinfo{author}{\bibfnamefont{S.~B.} \bibnamefont{Yuste}}, \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{M.}~\bibnamefont{{L\'opez de Haro}}},
  \bibinfo{journal}{Mol. Phys.} \textbf{\bibinfo{volume}{96}},
  \bibinfo{pages}{1} (\bibinfo{year}{1999}).

\bibitem[{\citenamefont{Santos et~al.}(2002)\citenamefont{Santos, Yuste, and
  {L\'opez de Haro}}}]{SYH02}
\bibinfo{author}{\bibfnamefont{A.}~\bibnamefont{Santos}},
  \bibinfo{author}{\bibfnamefont{S.~B.} \bibnamefont{Yuste}}, \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{M.}~\bibnamefont{{L\'opez de Haro}}},
  \bibinfo{journal}{J. Chem. Phys.} \textbf{\bibinfo{volume}{117}},
  \bibinfo{pages}{5785} (\bibinfo{year}{2002}).

\bibitem[{\citenamefont{{L\'opez de Haro} et~al.}(2008)\citenamefont{{L\'opez
  de Haro}, Yuste, and Santos}}]{HYS08}
\bibinfo{author}{\bibfnamefont{M.}~\bibnamefont{{L\'opez de Haro}}},
  \bibinfo{author}{\bibfnamefont{S.~B.} \bibnamefont{Yuste}}, \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{A.}~\bibnamefont{Santos}}, in
  \emph{\bibinfo{booktitle}{{Theory and Simulation of Hard-Sphere Fluids and
  Related Systems}}}, edited by
  \bibinfo{editor}{\bibfnamefont{A.}~\bibnamefont{Mulero}}
  (\bibinfo{publisher}{Springer}, \bibinfo{address}{Berlin},
  \bibinfo{year}{2008}), vol. \bibinfo{volume}{753} of
  \emph{\bibinfo{series}{Lectures Notes in Physics}}, pp.
  \bibinfo{pages}{183--245}.

\bibitem[{\citenamefont{Jenkins and Mancini}(1987)}]{JM87}
\bibinfo{author}{\bibfnamefont{J.~T.} \bibnamefont{Jenkins}} \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{F.}~\bibnamefont{Mancini}},
  \bibinfo{journal}{J. Appl. Mech.} \textbf{\bibinfo{volume}{54}},
  \bibinfo{pages}{27} (\bibinfo{year}{1987}).

\bibitem[{\citenamefont{Santos et~al.}(2005)\citenamefont{Santos, Yuste, and
  {L\'opez de Haro}}}]{SYH05}
\bibinfo{author}{\bibfnamefont{A.}~\bibnamefont{Santos}},
  \bibinfo{author}{\bibfnamefont{S.~B.} \bibnamefont{Yuste}}, \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{M.}~\bibnamefont{{L\'opez de Haro}}},
  \bibinfo{journal}{J. Chem. Phys.} \textbf{\bibinfo{volume}{123}},
  \bibinfo{pages}{234512} (\bibinfo{year}{2005}).

\bibitem[{\citenamefont{{L\'opez de Haro} et~al.}(2006)\citenamefont{{L\'opez
  de Haro}, Yuste, and Santos}}]{HYS06}
\bibinfo{author}{\bibfnamefont{M.}~\bibnamefont{{L\'opez de Haro}}},
  \bibinfo{author}{\bibfnamefont{S.~B.} \bibnamefont{Yuste}}, \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{A.}~\bibnamefont{Santos}},
  \bibinfo{journal}{Mol. Phys.} \textbf{\bibinfo{volume}{104}},
  \bibinfo{pages}{3461} (\bibinfo{year}{2006}).

\bibitem[{\citenamefont{Reiss et~al.}(1959)\citenamefont{Reiss, Frisch, and
  Lebowitz}}]{RFL59}
\bibinfo{author}{\bibfnamefont{H.}~\bibnamefont{Reiss}},
  \bibinfo{author}{\bibfnamefont{H.~L.} \bibnamefont{Frisch}},
  \bibnamefont{and} \bibinfo{author}{\bibfnamefont{J.~L.}
  \bibnamefont{Lebowitz}}, \bibinfo{journal}{J. Chem. Phys.}
  \textbf{\bibinfo{volume}{31}}, \bibinfo{pages}{369} (\bibinfo{year}{1959}).

\bibitem[{\citenamefont{Helfand et~al.}(1961)\citenamefont{Helfand, Frisch, and
  Lebowitz}}]{HFL61}
\bibinfo{author}{\bibfnamefont{E.}~\bibnamefont{Helfand}},
  \bibinfo{author}{\bibfnamefont{H.~L.} \bibnamefont{Frisch}},
  \bibnamefont{and} \bibinfo{author}{\bibfnamefont{J.~L.}
  \bibnamefont{Lebowitz}}, \bibinfo{journal}{J. Chem. Phys.}
  \textbf{\bibinfo{volume}{34}}, \bibinfo{pages}{1037} (\bibinfo{year}{1961}).

\bibitem[{\citenamefont{Lebowitz et~al.}(1965)\citenamefont{Lebowitz, Helfand,
  and Praestgaard}}]{LHP65}
\bibinfo{author}{\bibfnamefont{J.~L.} \bibnamefont{Lebowitz}},
  \bibinfo{author}{\bibfnamefont{E.}~\bibnamefont{Helfand}}, \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{E.}~\bibnamefont{Praestgaard}},
  \bibinfo{journal}{J. Chem. Phys.} \textbf{\bibinfo{volume}{43}},
  \bibinfo{pages}{774} (\bibinfo{year}{1965}).

\bibitem[{\citenamefont{Mandell and Reiss}(1975)}]{MR75}
\bibinfo{author}{\bibfnamefont{M.}~\bibnamefont{Mandell}} \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{H.}~\bibnamefont{Reiss}}, \bibinfo{journal}{J.
  Stat. Phys.} \textbf{\bibinfo{volume}{13}}, \bibinfo{pages}{113}
  (\bibinfo{year}{1975}).

\bibitem[{\citenamefont{Rosenfeld}(1988)}]{R88}
\bibinfo{author}{\bibfnamefont{Y.}~\bibnamefont{Rosenfeld}},
  \bibinfo{journal}{J. Chem. Phys.} \textbf{\bibinfo{volume}{89}},
  \bibinfo{pages}{4272} (\bibinfo{year}{1988}).

\bibitem[{\citenamefont{Heying and Corti}(2004)}]{HC04b}
\bibinfo{author}{\bibfnamefont{M.}~\bibnamefont{Heying}} \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{D.}~\bibnamefont{Corti}}, \bibinfo{journal}{J.
  Phys. Chem. B} \textbf{\bibinfo{volume}{108}}, \bibinfo{pages}{19756}
  (\bibinfo{year}{2004}).

\bibitem[{\citenamefont{Lebowitz}(1964)}]{L64}
\bibinfo{author}{\bibfnamefont{J.~L.} \bibnamefont{Lebowitz}},
  \bibinfo{journal}{Phys. Rev.} \textbf{\bibinfo{volume}{133}},
  \bibinfo{pages}{A895} (\bibinfo{year}{1964}).

\bibitem[{\citenamefont{Boubl\'{\i}k}(1970)}]{B70}
\bibinfo{author}{\bibfnamefont{T.}~\bibnamefont{Boubl\'{\i}k}},
  \bibinfo{journal}{J. Chem. Phys.} \textbf{\bibinfo{volume}{53}},
  \bibinfo{pages}{471} (\bibinfo{year}{1970}).

\bibitem[{\citenamefont{Mansoori et~al.}(1971)\citenamefont{Mansoori, Carnahan,
  Starling, and T.~W.~Leland}}]{MCSL71}
\bibinfo{author}{\bibfnamefont{G.~A.} \bibnamefont{Mansoori}},
  \bibinfo{author}{\bibfnamefont{N.~F.} \bibnamefont{Carnahan}},
  \bibinfo{author}{\bibfnamefont{K.~E.} \bibnamefont{Starling}},
  \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{J.}~\bibnamefont{T.~W.~Leland}},
  \bibinfo{journal}{J. Chem. Phys.} \textbf{\bibinfo{volume}{54}},
  \bibinfo{pages}{1523} (\bibinfo{year}{1971}).

\bibitem[{\citenamefont{Santos et~al.}(2011)\citenamefont{Santos, Yuste, and
  {L\'opez de Haro}}}]{SYH11}
\bibinfo{author}{\bibfnamefont{A.}~\bibnamefont{Santos}},
  \bibinfo{author}{\bibfnamefont{S.~B.} \bibnamefont{Yuste}}, \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{M.}~\bibnamefont{{L\'opez de Haro}}},
  \bibinfo{journal}{J. Chem. Phys.} \textbf{\bibinfo{volume}{135}},
  \bibinfo{pages}{181102} (\bibinfo{year}{2011}).

\end{thebibliography}

\documentclass[reprint,twocolumn,amsmath,amssymb,aps,prb]{revtex4}

\usepackage{graphicx}% Include figure files
\usepackage{dcolumn}% Align table columns on decimal point
\usepackage{bm}% bold math

 \usepackage{color}

\usepackage{amssymb}
\usepackage{graphicx}
\usepackage{amsbsy}

\newcommand\beq{\begin{equation}}
\newcommand\eeq{\end{equation}}
\newcommand\beqa{\begin{eqnarray}}
\newcommand\eeqa{\end{eqnarray}}
\newcommand{\nn}{\nonumber\\}

\newcommand{\NN}{s}

\newcommand{\dd}{\text{d}}
\newcommand{\ex}{\text{ex}}

\newcommand{\pure}{\text{pure}}

\begin{document}

% Use the \preprint command to place your local institutional report
% number in the upper righthand corner of the title page in preprint mode.
% Multiple \preprint commands are allowed.
% Use the 'preprintnumbers' class option to override journal defaults
% to display numbers if necessary
%\preprint{}

%Title of paper

\title{Note: An exact scaling relation for truncatable free energies of polydisperse hard-sphere mixtures}

% repeat the \author .. \affiliation  etc. as needed
% \email, \thanks, \homepage, \altaffiliation all apply to the current
% author. Explanatory text should go in the []'s, actual e-mail
% address or url should go in the {}'s for \email and \homepage.
% Please use the appropriate macro foreach each type of information

% \affiliation command applies to all authors since the last
% \affiliation command. The \affiliation command should follow the
% other information
% \affiliation can be followed by \email, \homepage, \thanks as well.
\author{Andr\'es Santos}

\email{andres@unex.es}
\homepage{http://www.unex.es/eweb/fisteor/andres/}

\affiliation{Departamento de F\'{\i}sica, Universidad de
Extremadura, Badajoz, E-06071, Spain}

\date{\today}

\begin{abstract}
A theoretical model for polydisperse systems of hard spheres is said to be truncatable  when the excess free energy depends on the size distribution through a finite number $K$ of moments. This Note proves an exact scaling relation for truncatable free energies, which allows to reduce the effective degrees of freedom to $K-2$ independent combinations of the moments.
\end{abstract}
%Collaboration name if desired (requires use of superscriptaddress
%option in \documentclass). \noaffiliation is required (may also be
%used with the \author command).
%\collaboration can be followed by \email, \homepage, \thanks as well.
%\collaboration{}
%\noaffiliation

\date{\today}

% insert suggested keywords - APS authors don't need to do this
%\keywords{}

%\maketitle must follow title, authors, abstract, \pacs, and \keywords
\maketitle

Let us consider a polydisperse  hard-sphere  fluid mixture in $d$ dimensions with number density $\rho=N/V$, where $N$ is the total number of particles contained in a volume $V$. The number of particles with a diameter between $\sigma$ and $\sigma+\dd\sigma$ is $N x(\sigma)\dd\sigma$, so that $\int_0^\infty\dd\sigma \,x(\sigma)=1$. The  moments of the size distribution are
\beq
M_n\equiv \int_0^\infty\dd\sigma \,x(\sigma)\sigma^n.
\label{mom}
\eeq
The particular case of a discrete $\NN$-component mixture is included as $x(\sigma)=\sum_{i=1}^\NN x_i\delta(\sigma-\sigma_i)$ with $\sum_{i=1}^\NN x_i=1$.

Let $a^\ex[\rho,x(\sigma)]$  be the excess free energy per particle of the system in units of $k_BT$, where $k_B$ is the Boltzmann constant and $T$ is the absolute temperature. It is given by
\beq
a^\ex[\rho,x(\sigma)]=-N^{-1}\ln Q_N(V),
\label{aex}
\eeq
where
\beq
Q_N(V)=V^{-N}\int \dd\mathbf{r}_1\ldots\int \dd\mathbf{r}_N\prod_{1\leq i<j\leq N}\Theta\left(r_{ij}-\frac{\sigma_i+\sigma_j}{2}\right)
\label{QN}
\eeq
is the configuration integral, $\Theta(x)$ being the Heaviside step function.

We now  suppose that the mixture consists of a finite number $N_0=x_0 N$ of point particles (i.e., $\sigma_i=0$ for $1\leq i\leq N_0$) plus $N'=N-N_0$ particles ($N_0+1\leq j\leq N$) with a certain (continuous or discrete) size distribution $x'(\sigma)$. Thus,
\beq
x(\sigma)=x_0\delta(\sigma)+(1-x_0)x'(\sigma).
\label{x0}
\eeq
The number density of the ``bare'' mixture is $\rho'=N'/V=(1-x_0)\rho$ and its size moments  are related to those of the original composite mixture by
\beq
M_n'=\int_0^\infty\dd\sigma \,x'(\sigma)\sigma^n=\frac{M_n}{1-x_0},\quad n\geq 1.
\label{Mn}
\eeq
Setting $\sigma_i=0$  ($1\leq i\leq N_0$) in Eq.\ \eqref{QN} we get
\beq
Q_N(V)=(1-V'/V)^{N_0}Q_{N'}(V),
\label{QN'}
\eeq
where $V'=v_dN'M_d'=v_d N M_d$ is the volume excluded by the  $N'$ normal particles, $v_{d}=(\pi /4)^{d/2}/\Gamma (1+d/2)$ being the volume of a
$d$-dimensional sphere of unit diameter, and $Q_{N'}(V)$ is the configuration integral of the bare mixture.
In the derivation of Eq.\ \eqref{QN'} one first integrates over the
positions of the point particles, obtaining $(V-V')^{N'}$; the remaining integration over the normal
particles is then the same as in the system without point particles.
Therefore, in the special case \eqref{x0}, Eq.\ \eqref{aex} becomes
\beq
a^\ex[\rho,x(\sigma)]=-{x_0}\ln\left(1-v_d\rho M_d\right)+(1-x_0)a^\ex[\rho',x'(\sigma)].
\label{11b}
\eeq

Let us now go back to the general polydisperse case and assume a model free energy with ``truncatable'' structure,\cite{GKM82,SWC01,S02} i.e., the excess free energy $a^\ex[\rho,x(\sigma)]$ depends on the size distribution $x(\sigma)$ only through  a \emph{finite} number $K$ of moments $\{M_1, M_2,\ldots, M_K\}$. Dimensional analysis requires that the dependence  of $a^{\text{ex}}$ on $\rho$ and $\{M_1, M_2,\ldots, M_K\}$ takes place through the dimensionless combinations $\eta$ and $\{m_2,\ldots, m_K\}$, where $\eta\equiv v_d\rho M_d$ is the packing fraction and $m_n\equiv M_n/M_1^n$ are rescaled moments. Therefore, the truncatability hypothesis can be written as
\beq
a^\ex[\rho,x(\sigma)]=a^\ex(\eta;m_2,\ldots,m_K).
\label{1}
\eeq
It is important to bear in mind that Eq.\ \eqref{1} is not a rigorous property. For instance, Blaak\cite{B98} has exactly evaluated the fourth virial coefficient $B_4(\sigma_i,\sigma_j,\sigma_k,\sigma_\ell)$ for $d=3$ when the four diameters are such that the smallest sphere (say $\ell$) fits in the inner space made by the other three spheres being tangent to each other. The analytic expression of $B_4(\sigma_i,\sigma_j,\sigma_k,\sigma_\ell)$ is given as a linear combination of terms of the form $\sigma_i^{q_1}\sigma_j^{q_2}\sigma_k^{q_3}\sigma_\ell^{q_4}$ with $q_1+q_2+q_3+q_4=9$. However, while $q_1$, $q_2$, and $q_3$ are always not larger than $3$, terms with up to $q_4=9$ are present. This shows that the exact $B_4$ is, in contrast to what claimed in other works,\cite{B99b} incompatible with Eq.\ \eqref{1}, unless $K\geq 9$. Notwithstanding this, the truncatability hypothesis \eqref{1} is crucial from a practical point of view to reduce from functional to algebraic the phase transition problem  in polydisperse systems.\cite{SWC01} Moreover,
this hypothesis (with $K=3$) has recently received numerical support from simulation data of three-dimensional polydisperse mixtures,\cite{OL12} even for metastable states.

The objective now is to prove that, once the ansatz \eqref{1} is assumed, the exact relationship \eqref{11b} imposes a constraint under the form of a scaling law for $a^\ex(\eta;m_2,\ldots,m_K)$. First, note that, in the case of Eq.\ \eqref{x0}, Eq.\ \eqref{Mn} gives
\beq
m_n'=(1-x_0)^{n-1}{m_n},\quad n\geq 2.
\label{12}
\eeq
Next, assuming Eq.\ \eqref{1} and making use of Eq.\ \eqref{11b} one obtains
\begin{widetext}
  \beq
a^\ex(\eta;m_2,m_3,\ldots,m_K)+\ln(1-\eta)=\lambda\left[a^\ex(\eta;\lambda m_2,\lambda^2m_3,\ldots,\lambda^{K-1}m_K)+\ln(1-\eta)\right],
\label{scal}
\eeq
where $\lambda= 1-x_0$. The \emph{scaling} property \eqref{scal} is the main result of this Note. It implies the equivalent form
\beq
a^\ex(\eta;m_2,m_3,\ldots,m_K)
=
-\ln(1-\eta)+\frac{1}{m_2}
\mathcal{A}\left(\eta; \frac{m_3}{m_2^2},\ldots,\frac{m_K}{m_2^{K-1}}\right),
\label{A}
\eeq
\end{widetext}
where the scaling function $\mathcal{A}$ remains undetermined.
In the special case of a one-component system, $m_n=1$ and thus
\beq
\mathcal{A}(\eta;1,\ldots,1)=a^\ex_\pure(\eta)+\ln(1-\eta),
\label{pure}
\eeq
where $a^\ex_\pure(\eta)$ is the excess free energy  of the pure fluid. Scaling relations similar to \eqref{scal} and \eqref{A} can be obtained for the compressibility factor  $Z\equiv p/\rho k_BT$, where $p$ is the pressure,   by the thermodynamic relation $Z=1+{\eta}{\partial a^\ex}/{\partial\eta}$.

Equation  \eqref{A} constrains possible theoretical models with truncatable structure. In fact, it represents a significant reduction in the number of ``degrees of freedom'' of the excess free energy of a polydisperse HS fluid. The exact free energy $a^\ex$ is a \emph{functional} of the size distribution $x(\sigma)$ and thus it has an infinite number of degrees of freedom. The truncatability hypothesis reduces the number of independent variables to the first $K$ moments  $\{M_n, n=1,\ldots,K\}$, apart from the number density $\rho$. Dimensional analysis  trivially reduces that number to $K-1$ dimensionless moments $\{m_n, n=2,\ldots,K\}$, as indicated in Eq.\ \eqref{1}. Further, Eq.\ \eqref{A} makes explicit the dependence on $m_2$, thus reducing the effective number of independent quantities to $K-2$ ratios $m_n/m_2^{n-1}$ ($n=3,\ldots,K$).

The exact excess free energy in the one-dimensional case ($d=1$) is $a^\ex=-\ln(1-\eta)$ and thus it trivially verifies Eq.\ \eqref{A} with $\mathcal{A}=0$. The exact result for hard disks ($d=2$) is not known. However, assuming truncatability  with $K=2$, Eqs.\ \eqref{A} and \eqref{pure} imply
\beq
a^\ex(\eta;m_2)=-\ln(1-\eta)+\frac{1}{m_2}\left[a_\pure^\ex(\eta)+\ln(1-\eta)\right].
\label{K2}
\eeq
This result agrees with that derived by independent methods\cite{SYH99,SYH02,HYS08} and includes Jenkins and Mancini's equation of state\cite{JM87} as a particular case.
Finally, let us consider the three-dimensional case ($d=3$) with $K=3$. Now the scaling function $\mathcal{A}(\eta;y)$ depends on the ratio $y=m_3/m_2^2$ but otherwise it is arbitrary. Let us assume the explicit functional form $\mathcal{A}(\eta;y)=y^{-2}\left[\mathcal{A}_0(\eta)+\mathcal{A}_1(\eta)y\right]$.   Equation \eqref{pure} then implies that $\mathcal{A}_0(\eta)=a^\ex_\pure(\eta)+\ln(1-\eta)-\mathcal{A}_1(\eta)$. Therefore, Eq.\ \eqref{A} yields
\beqa
a^\ex(\eta;m_2,m_3)&=&-\ln(1-\eta)+\frac{m_2^3}{m_3^2}\left[a_\pure^\ex(\eta)+\ln(1-\eta)\right]\nn
&&+
\left(\frac{m_2}{m_3}-\frac{m_2^3}{m_3^2}\right)\mathcal{A}_1(\eta),
\label{K3}
\eeqa
where the function $\mathcal{A}_1(\eta)$ is not constrained by Eq.\ \eqref{pure}. Equation \eqref{K3} with $\mathcal{A}_1(\eta)=3\eta/(1-\eta)$ has been derived by a different route\cite{SYH05,HYS06,HYS08} and includes, as particular cases, the Scaled Particle
Theory,\cite{RFL59,HFL61,LHP65,MR75,R88,HC04b} the virial Percus--Yevick,\cite{L64} and the Boubl\'{\i}k--Mansoori--Carnahan--Starling--Leland\cite{B70,MCSL71} equations of state.
More generally, Eq.\ \eqref{K3} with free $\mathcal{A}_1(\eta)$ or Eq.\ \eqref{A} with $K=3$ and free $\mathcal{A}(\eta,y)$ can be useful to infer the equation of state of the pure system in the metastable region from measurements made on multi-component systems.\cite{SYH11}

The author is grateful to two anonymous reviewers for their constructive comments. Financial support from the Spanish Government through Grant No. FIS2010-16587 and from the Junta de Extremadura (Spain) through Grant No.\ GR10158 (partially financed by FEDER funds) is  acknowledged.

\bibliographystyle{apsrev}
\bibliography{D:/Dropbox/Public/bib_files/liquid}
\end{document}